OpenStudy (anonymous):

What is the degree of the polynomial -x^2(x-3)(x+2)(x-5)^2?

OpenStudy (anonymous):

I thought the highest power was the degree (2, in this case), but the answer it's giving me is 6. How is it 6?

OpenStudy (anonymous):

yes the answer is 6

OpenStudy (anonymous):

u can take it to simpliest form and see....... which will take time only......

OpenStudy (anonymous):

so I have got a faster way to do it

OpenStudy (anonymous):

the highest power of x in (x-5)^2 will be two. right?

OpenStudy (anonymous):

With you so far.

OpenStudy (anonymous):

similarly the highest power of x in (x-3)and (x+2) will be one. right?

OpenStudy (anonymous):

I'm guessing we are going to be adding all the powers up?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so u will get 2+1+1+2=6

OpenStudy (anonymous):

which will be the highest power of x

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Why can you do that here though? Like take -9x^3+9x-6. You couldn't add up the powers of 3 and 1 to get a degree of 4, could you?

OpenStudy (anonymous):

no the case here is different.....

OpenStudy (anonymous):

It is like this x^2 * x * x * x^2 = x^(2+1+1+2) = x^6

OpenStudy (anonymous):

Ah, Ok, I think I got it now. Just muliplying them out to make it simpler kinda. Well thanks, saura.

OpenStudy (anonymous):

yeah